OpenStudy (anonymous):

If charlie and nick work together, they can paint the house in 20 hours. harlie can do the job alone in 36 hours. How long would it take nick to do the job?

OpenStudy (whpalmer4):

Find the rates at which they do the work together and separately. Charlie takes 36 hours to paint 1 hours, so he paints at a rate of 1 house/36 hours or (1/36). Charlie and Nick together paint 1 hours/20 hours, or (1/20). \[\frac{\text{work}}{\text{rate}} = \text{time}\] \[\frac{1}{(\frac{1}{36}+\frac{1}{N})} = 20\]where \(N\) is the number of hours that Nick takes to paint the house by himself.

OpenStudy (anonymous):

how come it's 1 hours ?

OpenStudy (anonymous):

oh nvm i get it, thanks <3

OpenStudy (whpalmer4):

1 job — but we would set the problem up similarly if we were working a problem where we had actual quantities to work with. I was in a bit of a hurry and didn't stick the units in the equation, but it would have been better had I done so. Glad you figured it out!